OpenStudy (anonymous):

There are three jars, 2/9th of each of which is filled with pickle. If all this was to poured out to fill new jars (of the same size and shape) completely, how many jars could be filled?

OpenStudy (anonymous):

2/9th filled bye one jar so we have to fill until full jar is filled that is let x if x jar is filled than x * 2/9 = 1 that means x = 9/2 = 4.5 jar

OpenStudy (mathmate):